### v1

# Revisiting Multi-Agent Asynchronous Online Optimization with Delays: the Strongly Convex Case

Abstract
We revisit multi-agent asynchronous online optimization with delays, where only one of the agents becomes active for making the decision at each round, and the corresponding feedback is received by all the agents after unknown delays. Although previous studies have established an $O(\sqrt{dT})$ regret bound for this problem, they assume that the maximum delay $d$ is knowable or the arrival order of feedback satisfies a special property, which may not hold in practice. In this paper, we surprisingly find that when the loss functions are strongly convex, these assumptions can be eliminated, and the existing regret bound can be significantly improved to $O(d\log T)$ meanwhile. Specifically, to exploit the strong convexity of functions, we first propose a delayed variant of the classical follow-the-leader algorithm, namely FTDL, which is very simple but requires the full information of functions as feedback. Moreover, to handle the more general case with only the gradient feedback, we develop an approximate variant of FTDL by combining it with surrogate loss functions. Experimental results show that the approximate FTDL outperforms the existing algorithm in the strongly convex case.
1 Introduction
Online learning (Shalev-Shwartz, 2011; Hazan, 2016) has received great attention over the past decades, due to the emergence of large-scale applications such as online advertising (Tan and Srikant, 2012), recommendation systems (Lin et al., 2020), and portfolio management (Agarwal et al., 2006). In general, it is formulated as a repeated game between an agent and an adversary, where the agent first selects a decision $\mathbf{x}_{t}$ from a set $\mathcal{K}\subseteq\mathbb{R}^{n}$ at each round $t$ and then the adversary selects a loss function $f_{t}(\cdot):\mathcal{K}\mapsto\mathbb{R}$. The performance of the agent is measured by regret
$R_{T}=\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$
which is the gap between the cumulative loss of the agent and that of a fixed optimal decision over the total $T$ rounds. The most popular paradigm for online learning is online convex optimization (OCO) (Zinkevich, 2003; Hazan et al., 2007; Abernethy et al., 2008; Duchi et al., 2011; Kingma and Ba, 2015; Wang et al., 2020; Wan and Zhang, 2021, 2022), which focuses on the case with convex loss functions and decision sets. By exploiting the general convexity, these previous studies on OCO have achieved an optimal regret bound of $O(\sqrt{T})$. Moreover, if the loss functions are strongly convex, an optimal regret bound of $O(\log T)$ can be further achieved.
However, the standard OCO assumes that the feedback about the loss function $f_{t}(\cdot)$ can be received immediately after the decision $\mathbf{x}_{t}$ is made, which may not hold in many practical applications, especially those with distributed scenarios. For example, in distributed sensor networks (Olfati-Saber, 2007; Tsagkatakis and Savakis, 2011; Kia et al., 2015; Souza et al., 2016), multiple sensors capable of collecting, processing, and exchanging data are widely used in online detection, classification, and tracking of some targets. Given the limited capabilities of sensors, fully decentralized communication and computation are crucial to avoid overloading small sensors or overwhelming busy networks. For the same reason, only the sensor located near the target is activated to perform detection or classification at a specific moment, and communicates with its neighboring sensors for collaboration. Note that both the processes of target observation and data communication between these sensors will cause delays. Moreover, due to the possible heterogeneity of these sensors, different sensors collaborate asynchronously and suffer different delays in general.
Motivated by these applications, Hsieh et al. (2022) extend the standard OCO into multi-agent asynchronous online optimization with delays. The setting involves a set of agents =$\mathcal{M}=\{1,\ldots,M\}$, and only one agent becomes active in each round $t$ to make the decision $\mathbf{x}_{t}$. Due to the effect of delays and asynchronicities, each agent $i\in\mathcal{M}$ can only receive the feedback about $f_{t}(\cdot)$ at the end of round $t+d_{t,i}-1$, where $d_{t,i}\in\mathbb{Z}^{+}$ denotes an arbitrary delay. It is well-known that if only one agent exists, i.e., $M=1$, various algorithms have been proposed to achieve an optimal $O(\sqrt{dT})$ regret bound for convex functions and an optimal $O(d\log T)$ regret bound for strongly convex functions in the worst case (Weinberger and Ordentlich, 2002; Joulani et al., 2013; Quanrud and Khashabi, 2015; Joulani et al., 2016; Wan et al., 2022b), where $d=\max_{t,i}\{d_{t,i}\}$ denotes the maximum delay. However, compared with the single-agent case, multi-agent OCO with delays faces two new challenges: the non-monotonicity of feedback sequence and lack of synchronization, which make the time-decaying and even delay-dependent learning rates utilized in these single-agent delayed algorithms unavailable (Hsieh et al., 2022).
To tackle these challenges, Hsieh et al. (2022) propose the first algorithm called delayed dual averaging (DDA) for the multi-agent case. If the maximum delay $d$ is knowable or the arrival order of feedback satisfies a special property, they prove that DDA can utilize only local information to adjust the learning rate and achieve the optimal $O(\sqrt{dT})$ regret bound for convex functions. However, these assumptions may also not hold in practice, and thus it is natural to ask whether these assumptions can be removed. Moreover, it remains unclear whether the strong convexity of functions can be further utilized to improve the regret as in the single-agent case. In this paper, we provide an affirmative answer to both questions by proposing new algorithms that can exploit the strong convexity of functions to reduce the regret bound without any prior information and assumption about the delay. Our key idea is to extend the classical follow-the-leader (FTL) algorithm (Hannan, 1957; Kalai and Vempala, 2005; Hazan et al., 2007), which originally is parameter-free and enjoys an $O(\log T)$ regret bound for the standard OCO with strongly convex functions, into the multi-agent case. Specifically, we first consider the case with the delayed full-information feedback, and propose follow-the-delayed-leader (FTDL) by replacing all historical loss functions minimized in each round of FTL with all received loss functions. Our analysis reveals that FTDL can enjoy an $O(d\log T)$ regret bound for strongly convex functions in a parameter-free manner, similar to FTL. Furthermore, we study the more general case with only delayed gradient feedback, and further develop an approximate variant of our FTDL by combining it with gradient-based surrogate loss functions (Hazan et al., 2007). It is worth noting that this approximate variant not only enjoys the same $O(d\log T)$ regret bound for strongly convex functions, but can be even implemented more efficiently than FTDL. Finally, we conduct experiments on four publicly available datasets to demonstrate that our approximate FTDL can outperform the existing DDA algorithm in the case with strongly convex functions.
2 Related Work
In this section, we briefly review the related work about the standard OCO, and multi-agent OCO with delays.
2.1 The Standard OCO
Since the pioneering work of Zinkevich (2003), there has been a growing research interest in developing and analyzing algorithms for OCO (Hazan et al., 2007; Abernethy et al., 2008; Duchi et al., 2011; Kingma and Ba, 2015; Wang et al., 2020; Wan and Zhang, 2021, 2022). Roughly speaking, these existing OCO algorithms can be divided into two categories. The first category includes online gradient descent (OGD) (Zinkevich, 2003) and its variants, which can be viewed as an extension from gradient descent in offline optimization (Boyd and Vandenberghe, 2004) into the online setting. The other category contains follow-the-leader (FTL) (Hannan, 1957) and its variants, where the so-called “leader” is the optimal decision for all historical loss functions. Note that both categories are able to achieve an $O(\sqrt{T})$ regret bound for convex functions, and an $O(\log T)$ regret bound for strongly convex functions (Zinkevich, 2003; Hazan et al., 2007; Hazan, 2016). Moreover, these $O(\sqrt{T})$ and $O(\log T)$ regret bounds have been proved to be optimal for convex and strongly convex functions, respectively (Abernethy et al., 2008).
Here, we only introduce the detailed procedures about FTL and its variants, because they are more related to this paper. Specifically, at each round $t$, FTL simply chooses the following decision
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t-1}f_{s}(\mathbf{x}).$ (1)
Despite of the simplicity, it can enjoy the optimal $O(\log T)$ regret bound for strongly convex functions (Hazan et al., 2007). However, there are two unsatisfactory points that limit the application of FTL. First, one can prove that in the case with only general convex functions, the regret of FTL could be $\Omega(T)$ (Hazan, 2016). Second, to implement FTL, we need to solve an offline optimization problem per round, which may become a computational bottleneck in practice.
To address these two issues simultaneously, previous studies (Hannan, 1957; Kalai and Vempala, 2005; Shalev-Shwartz and Singer, 2006; Hazan, 2016) have proposed to apply FTL with a linearized approximation of the original loss functions, i.e.,
$\tilde{f}_{t}(\mathbf{x})=f_{t}(\mathbf{x}_{t})+\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{x}-\mathbf{x}_{t}\rangle$
and a randomized or strongly convex regularization $\mathcal{R}(\cdot):\mathcal{K}\mapsto\mathbb{R}$. In this way, the decision for each round $t$ becomes
$\begin{split}\mathbf{x}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t-1}\langle\nabla f_{s}(\mathbf{x}_{s}),\mathbf{x}\rangle+\frac{\mathcal{R}(\mathbf{x})}{\eta_{t}}\end{split}$
where $\eta_{t}$ is a parameter for trading off the approximate losses and regularization. Note that when $\mathcal{R}(\cdot)$ is chosen to be $1$-strongly convex, e.g., $\mathcal{R}(\mathbf{x})=\|\mathbf{x}\|_{2}^{2}/2$, this variant of FTL is called follow-the-regularized-leader (FTRL) or dual averaging, and can achieve the optimal $O(\sqrt{T})$ regret bound for convex functions by setting an appropriate $\eta_{t}$ (Shalev-Shwartz and Singer, 2006; Hazan, 2016).
Moreover, Hazan et al. (2007) consider $\lambda$-strongly convex functions, and propose to apply FTL with the following approximation
$\tilde{f}_{t}(\mathbf{x})=f_{t}(\mathbf{x}_{t})+\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{x}-\mathbf{x}_{t}\rangle+\frac{\lambda}{2}\|\mathbf{x}-\mathbf{x}_{t}\|_{2}^{2}.$ (2)
The corresponding decision for each round $t$ can be written as follows
$\begin{split}\mathbf{x}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t-1}\left(\langle\nabla f_{s}(\mathbf{x}_{s}),\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}-\mathbf{x}_{s}\|_{2}^{2}\right).\end{split}$
This variant of FTL is called follow-the-approximate-leader (FTAL) and can enjoy the optimal $O(\log T)$ regret bound for strongly convex functions (Hazan et al., 2007).
2.2 Multi-Agent OCO with Delays
The study of OCO with delays can date back to Weinberger and Ordentlich (2002), who consider the case with only one agent and a fixed delay, i.e., $d_{t}=d$ for any $t\in[T]$. They propose a black-box technique that maintains $d$ instances of a standard OCO algorithm and alternately utilize these instances to generate the new decision. In this way, each instance is utilized once every $d$ rounds, and actually only needs to handle a subsequence of the total $T$ rounds as in the non-delayed case. By combining this black-box technique with the optimal algorithms for OCO, one can achieve $O(\sqrt{dT})$ and $O(d\log T)$ regret bounds for convex and strongly convex functions, respectively. Note that as proved by Weinberger and Ordentlich (2002), these delay-dependent regret bounds are optimal in the worst case. Later, Joulani et al. (2013) extend the black-box technique of Weinberger and Ordentlich (2002) into the case with one agent and arbitrary delays. Even without prior information of delays, they achieve the same regret bounds for convex and strongly convex functions. However, these black-box techniques (Weinberger and Ordentlich, 2002; Joulani et al., 2013) require much higher memory costs than the standard OCO algorithm. To address this limitation, there has been growing research interest in developing memory-efficient delayed variants of various standard OCO algorithms (Langford et al., 2009; McMahan and Streeter, 2014; Quanrud and Khashabi, 2015; Joulani et al., 2016; Li et al., 2019; Héliou et al., 2020; Flaspohler et al., 2021; Wan et al., 2022a, b, 2024a, 2024b; Wu et al., 2024).
Despite the great flourish of research on delayed OCO, the general case with multiple agents has rarely been investigated. As emphasized by Hsieh et al. (2022), multi-agent OCO with delays is more challenging due to the following two reasons.
  • First, these agents may suffer different communication latency, and thus can only receive the delayed feedback in an asynchronous way. Due to the asynchronicity and change of the active agent, the amount of available feedback is generally no longer monotone.
  • Second, to promote high privacy and information security, it is common to assume that agents do not have access to a global counter that indicates how many decisions have been made at any given stage, and any non-local information.
To tackle these challenges, Hsieh et al. (2022) propose delayed dual averaging (DDA) that computes the decision for each round $t$ as follows
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\langle\nabla f_{s}(\mathbf{x}_{s}),\mathbf{x}\rangle+\frac{\mathcal{R}(\mathbf{x})}{\eta_{t}}$
where $\mathcal{F}_{t}$ is a set including the timestamps of all feedback received by the $t$-th active agent. They first prove that DDA can achieve the $O(\sqrt{dT})$ regret bound for convex functions, by setting
$\eta_{t}=O\left(1/\sqrt{d\left(|\mathcal{F}_{t}|+d\right)}\right).$
Unfortunately, the value of $d$ is generally unknown in practice, which could limit the application of DDA. [^1] Even with the prior information of $d$, Hsieh et al. (2022) actually also need to assume that $|\mathcal{F}_{s}|\leq|\mathcal{F}_{t}|$ for any $s\in\mathcal{F}_{t}$. Note that in the single agent case, this issue can be solved by utilizing the standard “doubling trick” (Cesa-Bianchi et al., 1997) to estimate the value of $d$ on the fly and then adaptively adjust the parameter $\eta$. However, it is highly non-trivial to extend this idea into the multi-agent case due to the lack of a global counter. As a result, Hsieh et al. (2022) propose an alternative mechanism inspired by data-dependent OCO algorithms, e.g., AdaGrad (Duchi et al., 2011), to adjust $\eta_{t}$ without prior information of $d$, but introduce an additional assumption on the arrival order of feedback. It remains unclear whether the regret of multi-agent OCO with delays can be bounded without any prior information and assumption about the delay.
3 Our Results
In this section, we first introduce some common assumptions, and then introduce our two algorithms as well as the corresponding theoretical guarantees.
3.1 Assumptions
Following previous studies (Shalev-Shwartz, 2011; Hazan, 2016; Hsieh et al., 2022), we first introduce two assumptions about the gradient norm of functions and the decision set.
Assumption 1.
The gradients of all loss functions are bounded by some constant $G$, i.e.,
$\|\nabla f_{t}(\mathbf{x})\|_{2}\leq G$
for any $\mathbf{x}\in\mathcal{K}$ and $t\in[T]$.
Assumption 2.
The decision set $\mathcal{K}$ contains the origin $\mathbf{0}$, and its diameter is bounded by some constant $D$, i.e.,
$\|\mathbf{x}-\mathbf{y}\|_{2}\leq D$
for any $\mathbf{x},\mathbf{y}\in\mathcal{K}$.
Moreover, as the critical difference between this paper and Hsieh et al. (2022), we introduce the following assumption about the strong convexity of functions.
Assumption 3.
All loss functions are $\lambda$-strongly convex over the decision set $\mathcal{K}$, i.e.,
$f_{t}(\mathbf{y})\geq f_{t}(\mathbf{x})+\langle\nabla f_{t}(\mathbf{x}),\mathbf{y}-\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}-\mathbf{y}\|^{2}_{2}$
for any $\mathbf{x},\mathbf{y}\in\mathcal{K}$ and $t\in[T]$.
3.2 Follow-the-Delayed-Leader
Before introducing our algorithms, we first provide a formal definition for the available information held by each agent $i\in[M]$. To be precise, let $\mathcal{G}_{i}$ denote the set including the timestamps for all feedback received by each agent $i\in[M]$. In the beginning, the set should be initialized as $\mathcal{G}_{i}=\emptyset$. Then, due to the effect of delays and asynchronicity, each agent $i$ will receive the feedback from some previous round $s$, and update the set as $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$. Moreover, it will relay the received feedback if necessary. In this way, when an agent $i$ is activated at round $t$, it can only make the decision based on the feedback from rounds in the latest set $\mathcal{G}_{i}$ that may only be a subset of $[t-1]$.
Algorithm 1 Follow-the-Delayed-Leader (Agent $i$)
1:  Initialize: $\mathcal{G}_{i}=\emptyset$ and $t=1$
2:  while not stopped do
3:     Asynchronously receive feedback $f_{s}(\mathbf{x})$ from round $s$
4:     $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$
5:     Relay $f_{s}(\mathbf{x})$ if necessary
6:     if the agent becomes active then
7:        $\mathcal{F}_{t}=\mathcal{G}_{i}$
8:        Play $\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}f_{s}(\mathbf{x})$
9:     end if
10:  end while
Therefore, even if the functions are strongly convex, FTL in (1) cannot be directly utilized. To address this issue, we propose a delayed variant of FTL for multi-agent OCO with delays that is called follow-the-delayed-leader (FTDL) and utilizes the following decision
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}f_{s}(\mathbf{x})$
where $\mathcal{F}_{t}$ is set to be the latest $\mathcal{G}_{i}$ of the active agent $i$ at round $t$. The detailed procedures of our FTDL are summarized in Algorithm 1 from the point of view of any single agent $i$. Despite of the simplicity, we prove FTDL enjoys the following regret bound, which demonstrates that it can exploit the strong convexity automatically.
Theorem 1.
Under Assumptions 1 and 3, Algorithm 1 ensures
$R_{T}\leq\frac{2dG^{2}(1+\ln T)}{\lambda}.$
Proof.
Inspired by the existing analysis for FTL (Garber and Hazan, 2016), we first define an ideal decision for each round $t=1,\dots,T$ as below
$\tilde{\mathbf{x}}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t}f_{s}(\mathbf{x})$
which actually is even one step ahead of FTL in (1). In the literature, it is well-known that such an ideal decision enjoys non-positive regret for any loss functions $f_{1}(\mathbf{x}),\dots,f_{T}(\mathbf{x})$ over $\mathcal{K}$.
Lemma 1.
(Lemma 6.6 in Garber and Hazan (2016)) Let $f_{1}(\mathbf{x}),\dots,f_{T}(\mathbf{x})$ be a sequence of loss functions over $\mathcal{K}$, and let $\tilde{\mathbf{x}}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{i=1}^{t}f_{i}(\mathbf{x})$, then
$\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\leq 0.$
Moreover, the regret of our FTDL can be divided into two parts: one is the regret of the ideal decision, and the other is the gap between the cumulative loss of the ideal decision and our FTDL, as follows
$R_{T}=$ $\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{x}^{\ast})$ (3)
[…]
where $\mathbf{x}^{\ast}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$.
By combining Lemma (1) with (3), it is easy to verify that
$R_{T}\leq$ $\sum_{t=1}^{T}\left(f_{t}({\mathbf{x}}_{t})-f_{t}(\tilde{\mathbf{x}}_{t})\right)\leq\sum_{t=1}^{T}\langle\nabla f_{t}({\mathbf{x}}_{t}),{\mathbf{x}}_{t}-\tilde{\mathbf{x}}_{t}\rangle$ (4)
$\leq$ $\sum_{t=1}^{T}\|\nabla f_{t}({\mathbf{x}}_{t})\|_{2}\|{\mathbf{x}}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}\leq G\sum_{t=1}^{T}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
where the last inequality is due to Assumption 1.
Then, we only need to bound the distance between our decision and the ideal decision. To this end, we introduce a nice property of strongly convex functions from Hazan and Kale (2012).
Lemma 2.
Assume $f(\mathbf{x})$ is $\lambda$-strongly convex over $\mathcal{K}$, it holds that
$\|\mathbf{x}-\mathbf{x}^{*}\|^{2}_{2}\leq\frac{2}{\lambda}\left((f(\mathbf{x})-f(\mathbf{x}^{*})\right)$
for $\mathbf{x}^{*}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}f(\mathbf{x})$ and any $\mathbf{x}\in\mathcal{K}$.
Let $F_{t}(\mathbf{x})=\sum_{s=1}^{t}f_{s}(\mathbf{x})$ for any $t=1,\dots,T$. Note that $F_{t}(\mathbf{x})$ is $(t\lambda)$-strongly convex over $\mathcal{K}$ due to Assumption 3. As a result, by combining Lemma 2 with the definition of $\tilde{\mathbf{x}}_{t}$ and $F_{t}(\mathbf{x})$, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|^{2}_{2}$
$\leq$ $\frac{2}{t\lambda}\left(F_{t}(\mathbf{x}_{t})-F_{t}(\tilde{\mathbf{x}}_{t})\right)$
$=$ $\frac{2}{t\lambda}\left(\sum_{s\in\mathcal{F}_{t}}f_{s}(\mathbf{x}_{t})-\sum_{s\in\mathcal{F}_{t}}f_{s}(\tilde{\mathbf{x}}_{t})\right)$
$+\frac{2}{t\lambda}\left(\sum_{s\in[t]\setminus\mathcal{F}_{t}}f_{s}(\mathbf{x}_{t})-\sum_{s\in[t]\setminus\mathcal{F}_{t}}f_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2}{t\lambda}\left(\sum_{s\in[t]\setminus\mathcal{F}_{t}}f_{s}(\mathbf{x}_{t})-\sum_{s\in[t]\setminus\mathcal{F}_{t}}f_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2}{t\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\|\nabla f_{s}(\mathbf{x}_{t})\|_{2}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
$\leq$ $\frac{2}{t\lambda}G\left(t-|\mathcal{F}_{t}|\right)\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
where the second inequality is due to the definition of $\mathbf{x}_{t}$.
By combining the above inequality with the fact that any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round $t-1$, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}\leq\frac{2G(t-|\mathcal{F}_{t}|)}{t\lambda}\leq\frac{2dG}{t\lambda}.$ (5)
Finally, by substituting (5) into (4), we have
$R_{T}\leq\frac{2dG^{2}}{\lambda}\sum_{t=1}^{T}\frac{1}{t}\leq\frac{2dG^{2}(1+\ln T)}{\lambda}.$
Remark.
First, from Theorem 1, our FTDL can exploit the strong convexity of functions to achieve an $O(d\log T)$ regret bound, which is much tighter than the $O(\sqrt{dT})$ regret bound achieved by the DDA algorithm (Hsieh et al., 2022). More interestingly, our FTDL does not require any prior information and assumption about the delay. Second, it is worth noting that the regret bound in Theorem 1 holds without Assumption 2, which implies that our FTDL can be directly applied to the unbounded set. The nice property actually inherits from the standard OCO with strongly convex functions (Hazan et al., 2007).
3.3 Approximate Follow-the-Delayed-Leader
Although our FTDL has already achieved the desired regret bound, two unsatisfactory points still exist, which could possibly limit its application. First, FTDL requires full-information feedback, i.e., the total information about each loss function $f_{t}(\mathbf{x})$. However, in general, one may only be able to receive the gradient of each loss function $f_{t}(\mathbf{x})$, i.e., $\nabla f_{t}(\mathbf{x}_{t})$. Second, similar to FTL, our FTDL also needs to solve an offline optimization problem to compute its decision, which may be time-consuming in practice. To address these issues, we further develop an approximate variant of FTDL, by combining it with gradient-based surrogate loss functions.
Specifically, inspired by the approximate variant of FTL (Hazan et al., 2007), it is natural to apply our FTDL to the surrogate loss functions defined in (2), which adopts the following decision
$\begin{split}\mathbf{x}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\nabla f_{s}(\mathbf{x}_{s}),\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}-\mathbf{x}_{s}\|_{2}^{2}\right).\end{split}$ (6)
At first glance, to compute $\mathbf{x}_{t}$ in (6), these agents need to share both the gradient $\nabla f_{t}(\mathbf{x}_{t})$ and the decision $\mathbf{x}_{t}$ with others, which doubles the communication costs of the DDA algorithm (Hsieh et al., 2022). However, it is worth noting that $\mathbf{x}_{t}$ in (6) is equivalent to
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\nabla f_{s}(\mathbf{x}_{s})-\lambda\mathbf{x}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}\right)$ (7)
because the term $(\lambda/2)\|\mathbf{x}_{s}\|_{2}^{2}$ does not affect the minimizer of the optimization problem in (6). As a result, these agents only need to share one vector $\mathbf{z}_{t}=\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}$ with others, which enjoys the same communication costs as the DDA algorithm (Hsieh et al., 2022).
Furthermore, we notice that when $\mathcal{F}_{t}=\emptyset$, i.e., the active agent at round $t$ have not received any feedback yet, the decision $\mathbf{x}_{t}$ generated by (7) could be arbitrary in $\mathcal{K}$. To minimize the randomness, we introduce an additional regularization into (7) as below
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}\right)+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$
which unifies $\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}=\mathbf{0}$ for the case with $\mathcal{F}_{t}=\emptyset$ under Assumption 2. The detailed procedures of our approximate FTDL are summarized in Algorithm 2 from the point of view of any single agent $i$.
Now, we demonstrate that it can also exploit the strong convexity of functions as the original FTDL.
Algorithm 2 Approximate FTDL (Agent $i$)
1:  Initialize: $\mathcal{G}_{i}=\emptyset$ and $t=1$
2:  while not stopped do
3:     Asynchronously receive feedback $\mathbf{z}_{s}=\nabla f_{s}(\mathbf{x}_{s})-\lambda\mathbf{x}_{s}$ from round $s$
4:     $\mathcal{G}_{i}=\mathcal{G}_{i}\cup\{s\}$
5:     Relay $\mathbf{z}_{s}$ if necessary
6:     if the agent becomes active then
7:        Set $\mathcal{F}_{t}=\mathcal{G}_{i}$ and play
$\mathbf{x}_{t}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s\in\mathcal{F}_{t}}\left(\langle\mathbf{z}_{s},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}\right)+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}$
8:     end if
9:  end while
Theorem 2.
Under Assumptions 1, 2, and 3, Algorithm 2 ensures
$R_{T}\leq\frac{\lambda D^{2}}{2}+\frac{2d(G+2\lambda D)^{2}(\ln T+1)}{\lambda}.$
Proof.
For any $t=1,\dots,T$, let $\tilde{f}_{t}(\mathbf{x})=\langle\mathbf{z}_{t},\mathbf{x}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$ and $F_{t}(\mathbf{x})=\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}$. Similar to the proof of Theorem 1, for each round $t$, we can define an ideal decision as below
$\begin{split}\tilde{\mathbf{x}}_{t}=&\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{s=1}^{t}\tilde{f}_{s}(\mathbf{x})+\frac{\lambda}{2}\|\mathbf{x}\|^{2}_{2}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}F_{t}(\mathbf{x}).\end{split}$ (8)
According to Assumption 3, we have
$R_{T}=$ $\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t})-\sum_{t=1}^{T}f_{t}(\mathbf{x}^{\ast})$ (9)
$\leq$ $\sum_{t=1}^{T}\left(\langle\nabla f_{t}(\mathbf{x}_{t}),\mathbf{x}_{t}-\mathbf{x}^{\ast}\rangle-\frac{\lambda}{2}\|\mathbf{x}_{t}-\mathbf{x}^{\ast}\|^{2}_{2}\right)$
$=$ $\sum_{t=1}^{T}\left(\tilde{f}_{t}(\mathbf{x}_{t})-\tilde{f}_{t}(\mathbf{x}^{\ast})\right)$
$=$ $\sum_{t=1}^{T}\left(\tilde{f}_{t}(\mathbf{x}_{t})-\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})\right)+\sum_{t=1}^{T}\left(\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})-\tilde{f}_{t}(\mathbf{x}^{\ast})\right)$
where $\mathbf{x}^{\ast}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$.
Then, we notice that the first term in the right side of (9) can be bounded by analyzing the distance of our decision and the ideal decision. Specifically, under Assumptions 1 and 2, for any $t=1,\dots,T$, we have
$\begin{split}&\|\mathbf{z}_{t}\|_{2}=\|\nabla f_{t}(\mathbf{x}_{t})-\lambda\mathbf{x}_{t}\|_{2}\\ \leq&\|\nabla f_{t}(\mathbf{x}_{t})\|_{2}+\|\lambda\mathbf{x}_{t}\|_{2}\leq G+\lambda D.\end{split}$ (10)
Moreover, for any $t=1,\dots,T$, we can show that $\tilde{f}(\mathbf{x})$ is $(G+2\lambda D)$-Lipschitz, i.e., for any $\mathbf{x},\mathbf{y}\in\mathcal{K}$, it holds that
$\begin{split}\left|\tilde{f}_{t}(\mathbf{x})-\tilde{f}_{t}(\mathbf{y})\right|=&\left|\langle\mathbf{z}_{t},\mathbf{x}-\mathbf{y}\rangle+\frac{\lambda}{2}\|\mathbf{x}\|_{2}^{2}-\frac{\lambda}{2}\|\mathbf{y}\|_{2}^{2}\right|\\ =&\left|\left\langle\mathbf{z}_{t}+\frac{\lambda}{2}(\mathbf{x}+\mathbf{y}),\mathbf{x}-\mathbf{y}\right\rangle\right|\\ \leq&\left\|\mathbf{z}_{t}+\frac{\lambda}{2}(\mathbf{x}+\mathbf{y})\right\|_{2}\|\mathbf{x}-\mathbf{y}\|_{2}\\ \leq&(G+2\lambda D)\|\mathbf{x}-\mathbf{y}\|_{2}\end{split}$ (11)
where the last inequality is due to (10) and Assumptions 1 and 2.
From (11), it is easy to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\mathbf{x}_{t})-\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})\right)\leq(G+2\lambda D)\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}.\end{split}$ (12)
Note that $F_{t}(\mathbf{x})$ is $((t+1)\lambda)$-strongly convex over $\mathcal{K}$. As a result, by combining Lemma 2 with the definition of $\tilde{\mathbf{x}}_{t}$ and $F_{t}(\mathbf{x})$, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|^{2}_{2}$
$\leq$ $\frac{2}{(t+1)\lambda}\left(F_{t}(\mathbf{x}_{t})-F_{t}(\tilde{\mathbf{x}}_{t})\right)$
$=$ $\frac{2}{(t+1)\lambda}\left(\sum_{s\in\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)+\frac{\lambda\left(\|\mathbf{x}_{t}\|_{2}^{2}-\|\tilde{\mathbf{x}}_{t}\|_{2}^{2}\right)}{2}\right)$
$+\frac{2}{(t+1)\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2}{(t+1)\lambda}\sum_{s\in[t]\setminus\mathcal{F}_{t}}\left(\tilde{f}_{s}(\mathbf{x}_{t})-\tilde{f}_{s}(\tilde{\mathbf{x}}_{t})\right)$
$\leq$ $\frac{2\left(t-|\mathcal{F}_{t}|\right)(G+2\lambda D)}{(t+1)\lambda}\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}$
where the second inequality is due to the definition of $\mathbf{x}_{t}$, and the last inequality is due to (11).
By combining the above inequality with the fact any feedback from rounds $1,\dots,t-d$ must have been received by any agent $i$ at the end of round $t-1$, we have
$\|\mathbf{x}_{t}-\tilde{\mathbf{x}}_{t}\|_{2}\leq\frac{2d(G+2\lambda D)}{(t+1)\lambda}.$ (13)
Now, we only need to bound the last term in the right side of (9). Specifically, by applying Lemma 1 with the definition of $\tilde{\mathbf{x}}_{t}$, we have
$\begin{split}\left(\sum_{t=1}^{T}\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})+\frac{\lambda}{2}\|\tilde{\mathbf{x}}_{0}\|_{2}^{2}\right)-\left(\sum_{t=1}^{T}\tilde{f}_{t}(\mathbf{x}^{\ast})+\frac{\lambda}{2}\|{\mathbf{x}}^{\ast}\|_{2}^{2}\right)\leq 0\end{split}$
where $\tilde{\mathbf{x}}_{0}=\operatorname*{argmin}_{\mathbf{x}\in\mathcal{K}}(\lambda/2)\|\mathbf{x}\|_{2}^{2}=\mathbf{0}$ under Assumption 1.
From the above inequality, it is not hard to verify that
$\begin{split}\sum_{t=1}^{T}\left(\tilde{f}_{t}(\tilde{\mathbf{x}}_{t})-\tilde{f}_{t}(\mathbf{x}^{\ast})\right)\leq\frac{\lambda}{2}\|\mathbf{x}^{\ast}\|_{2}^{2}\leq\frac{\lambda D^{2}}{2}\end{split}$ (14)
where the last inequality is also due to Assumption 1.
Finally, by combining (9) with (12), (13), and (14), we have
$\begin{split}R_{T}\leq&\frac{\lambda D^{2}}{2}+\sum_{t=1}^{T}\frac{2d(G+2\lambda D)^{2}}{(t+1)\lambda}\\ \leq&\frac{\lambda D^{2}}{2}+\frac{2d(G+2\lambda D)^{2}(\ln T+1)}{\lambda}.\end{split}$
Remark.
From Theorem 2, the approximate FTDL enjoys the same $O(d\log T)$ regret bound as FTDL, and also does not require any prior information or assumption about the delay. Moreover, we emphasize that although our Theorem 2 requires the boundedness of decision sets, i.e., Assumption 2, it is actually easy to extend our Algorithm 2 to further handle the unbounded set by using the fact that the optimal solution of a strongly convex function must lie in a ball centered at the origin (see Remark 2 of Wan et al. (2022b) for details).
(a) ijcnn1
(b) w8a
(c) phishing
(d) a9a
Figure 1: Comparisons of our A-FTDL against DDA on delayed $2$-agent online binary classification.
4 Experiments
In this section, we conduct simulation experiments on four publicly available datasets—ijcnn1, w8a, phishing and a9a from the LIBSVM repository (Chang and Lin, 2011) to compare the performance of our approximate FTDL (A-FTDL) against the DDA algorithm (Hsieh et al., 2022). The details of these datasets are summarized in Table 1. Both algorithms are implemented with Matlab and tested on a laptop with 2.6GHz CPU and 16GB memory.
Dataset # Examples ($T$) # Features ($n$) # Classes
ijcnn1 49990 22 2
w8a 49749 300 2
phishing 11055 68 2
a9a 32561 123 2
Table 1: Datasets used in our experiments.
Specifically, we consider a delayed $2$-agent online binary classification problem with strongly convex loss functions. For each dataset, we utilize all training examples to conduct our experiments, and thus $T$ is equal to the number of these training examples. In each round $t\in[T]$, one of the agent will be active, and choose a decision $\mathbf{x}_{t}$ from the following convex set
$\mathcal{K}=\{\mathbf{x}\in\mathbb{R}^{n}\mid\|\mathbf{x}\|_{2}\leq 1\}$
which satisfies Assumption  2 with $D=2$, where $n$ is equal to the number of features in the dataset. Then, an example $(\mathbf{w}_{t},y_{t})$ is selected from the dataset, and causes a loss
$f_{t}(\mathbf{x}_{t})=\max\left\{1-y_{t}\mathbf{w}_{t}^{\top}\mathbf{x}_{t},0\right\}+\frac{\lambda}{2}\|\mathbf{x}_{t}\|^{2}_{2}$
where $\mathbf{w}_{t}\in\mathbb{R}^{n}$ denotes the feature vector, $y_{t}\in\{-1,1\}$ denotes the class label, and $\lambda=0.01$. It is easy to verify that the loss function $f_{t}(\mathbf{x})$ satisfies Assumption 3, and ensures
$\|\nabla f_{t}(\mathbf{x})\|_{2}\leq\lambda\|\mathbf{x}\|_{2}+\|\mathbf{w}_{t}\|_{2}.$
Note that by a single pass of the datasets, we can compute the maximum norm of $\mathbf{w}_{t}$, and denote this value as $w_{max}$. Then, the above loss function $f_{t}(\mathbf{x})$ also satisfies Assumption 1 with
$G=\frac{\lambda D}{2}+w_{max}=0.01+w_{max}.$
Next, to simulate the delays, we randomly generate two delayed sequences for the two agents, each sequence is divided into four parts with delays selected from ranges of $[1,250]$, $[1,500]$, $[1,750]$, and $[1,1000]$, respectively. In this way, the maximum delay $d$ is equal to $1000$. Moreover, we recall that the DDA algorithm needs to choose an appropriate $\mathcal{R}(\mathbf{x})$ and tune the parameter $\eta_{t}$. According to Proposition 7 of Hsieh et al. (2022), we set $\mathcal{R}(\mathbf{x})=(1/2)\|\mathbf{x}\|_{2}^{2}$ and use
$\eta_{t}=\frac{D}{2\sqrt{2}G\sqrt{(1+2d)(|\mathcal{F}_{t}|+d+1)}}$
which ensures the $O(\sqrt{dT})$ regret bound.
Fig. 1 shows the comparisons of our A-FTDL and the DDA algorithm (Hsieh et al., 2022) over all these datasets. We find that the cumulative loss of A-FTDL is smaller than that of DDA, which verifies the theoretical improvement on the regret by exploiting the strong convexity of functions. Moreover, these results also demonstrate the benefit of strong convexity for removing the requirement on prior information and assumption about the delay.
5 Conclusion and Future Work
In this paper, we investigate multi-agent asynchronous online optimization with delays, and develop two novel algorithms called FTDL and A-FTDL for the case of strongly convex functions. According to the analysis, our two algorithms exploit the strong convexity to achieve an $O(d\log T)$ regret bound, which is much tighter than the existing $(\sqrt{dT})$ regret bound achieved by only using the convexity. More interestingly, different from the existing algorithm that needs to know the maximum delay beforehand or requires an additional assumption on the arrival order of feedback, our two algorithms do not require any prior information or assumption about the delay. Finally, numerical experiments show that our A-FTDL outperforms the existing algorithm in real applications.
An open question is whether our algorithms and analysis can be further extended to a broader class of exponentially concave (exp-concave) functions. We notice that in the standard OCO, there is also a variant of FTL (Hazan et al., 2007) that enjoys an $O(n\log T)$ regret bound for exp-concave functions. However, it seems highly non-trivial to exploit the exponential concavity in the delayed setting, even with only single agent (Wan et al., 2022b). Therefore, we leave this extension for future work.